OpenStudy (anonymous):

8. What is the equation of the line, in point-slope form, that passes through the given point and has the given slope? Point: (8,-8); Slope: 3 CAN SOMEONE EXPLAIN HOW TO DO THIS PLEASE A MEDAL IF SO??

OpenStudy (anonymous):

y-y1= m(x-x1) just substitute values in there, secondly its a medal which is given after choosing best response

OpenStudy (anonymous):

what's x and what y?

jimthompson5910 (jim_thompson5910):

the given point is (x1,y1), so because the given point is (8,-8), this means x1 = 8 y1 = 8 the given slope is m = 3

jimthompson5910 (jim_thompson5910):

so y-y1= m(x-x1) transforms into y-(-8)= 3(x-8) and that turns into y+8 = 3(x-8)

OpenStudy (anonymous):

its the standard format of a line and x and y are the points lying on the line

OpenStudy (anonymous):

Thank you very much !!